OpenStudy (anonymous):

@ myininaya where does the x+c come from in ln|y+1|=x+C p.s this was in reference to the question "if dy/dx = 1+y, and when x=0, y=2, show that y = 3e^x -1 " you had answered before

OpenStudy (anonymous):

I hope I'm not stepping on any toes here by submitting an answer, but in this separable differential equation, you have to integrate the dx term as well. When you integrate the dx, you get an x. You'll end up with a constant term C because you have an indefinite integral.

myininaya (myininaya):

jabberwock if you want to see what i did you can come here i think she/he is having a problem with it http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e1d1f2f0b8b4841c1a9a219 maybe you can give something helpful to the person because i guess im not explaining it well

OpenStudy (anonymous):

Nah, that looks better than how I would have explained it. Unless there's a specific question, I don't think I could do any better.

OpenStudy (anonymous):

LOL @ JABBERWOCK AND MYININAYA thank you math genuises!

OpenStudy (anonymous):

Lol

myininaya (myininaya):

np

OpenStudy (anonymous):

@ jabberwock, y+1=e^x*(3) how did e^{ln3} become (3)? and what happened to the absolute value brackets around y+1

OpenStudy (anonymous):

Raising e to a power and taking a natural log are inverse operations. This means that when you do both, they kind of cancel each other out, and you're left with whatever is on the inside. The absolute value is gone because e to any power is going to be positive. To verify, think about the graph of e^x. It is asymptotic to the line y=0.

OpenStudy (anonymous):

i don't get it =[

OpenStudy (anonymous):

myininaya, want to take a stab at it?

myininaya (myininaya):

ln(e^x)=x for all real x e^(lnx)=x for x>0

OpenStudy (anonymous):

wow i don't remember learning that

myininaya (myininaya):

what about when i did e^(x+C)=e^x*e^C or was it something else i can't remember exactly but it was something like that

OpenStudy (anonymous):

oh i get that part because that is just like x^ab = x^a +x^b

OpenStudy (anonymous):

3+a-a = 3 3-a+a = 3 3a/a = 3 (3/a)(a) = 3 \[\sqrt{3^2}=3\]\[(\sqrt{3})^2=3\] \[\ln(e^3)=3\]\[e^{\ln3}=3\] All these operations undo each other. In each case, you start with 3, do something to it, undo it, and get back to 3.

OpenStudy (anonymous):

OHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHH

OpenStudy (anonymous):

wow thank you so much, i don't think i had learnt that in the log rules. do you know where about in the topic it occurs or maybe i'm just a nutcase and skipped it completely

myininaya (myininaya):

x^(ab) does not equal x^a +x^b

OpenStudy (anonymous):

It's not explicitly taught a lot of times. It sometimes comes out in problem sets though.

myininaya (myininaya):

the inverse thing?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

so what does x^(ab) equal to? lol @jabberwork

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

\[(x^a)^b\]

OpenStudy (anonymous):

oh i see, lol i totally fail in math!

myininaya (myininaya):

lol

OpenStudy (anonymous):

thanks guys, i won't harass you no further! see you

OpenStudy (anonymous):

You're fine. Take care.